# Reinforced Diffuser for Red Teaming Large Vision-Language Models

### v1

@@
Abstract
The rapid advancement of large Vision-Language Models (VLMs) has raised significant safety concerns, particularly regarding their vulnerability to jailbreak attacks. While existing research primarily focuses on VLMs’ susceptibility to harmful instructions, this work identifies a critical yet overlooked vulnerability: current alignment mechanisms often fail to address the risks posed by toxic text continuation tasks. To investigate this issue, we propose a novel Red Team Diffuser (RTD) framework, which leverages reinforcement learning to generate red team images that effectively induce highly toxic continuations from target black-box VLMs. The RTD pipeline begins with a greedy search for high-quality image prompts that maximize the toxicity of VLM-generated sentence continuations, guided by a Large Language Model (LLM). These prompts are then used as input for the reinforcement fine-tuning of a diffusion model, which employs toxicity and alignment rewards to further amplify harmful outputs. Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% on the original attack set and 8.91% on a hold-out set. Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA. These findings reveal significant deficiencies in existing alignment strategies, particularly their inability to prevent harmful continuations. Our work underscores the urgent need for more robust and adaptive alignment mechanisms to ensure the safe deployment of VLMs in real-world applications.
@@ Abstract
footnotetext: $\dagger$  Corresponding author.
1 Introduction
The widespread deployment of large Vision-Language Models (VLMs), such as OpenAI’s GPT-4o Achiam et al. (2023) and Google’s Gemini Team et al. (2023), has raised significant concerns about their safety to potential jailbreak attacks. While VLMs inherit alignment properties from Large Language Models (LLMs), their multimodal nature introduces novel safety risks. Current jailbreak attacks on VLMs leverage either unimodal or multimodal approaches to manipulate the models into responding to harmful requests, such as providing bomb-making instructions Wang et al. (2024a).
However, research remains limited on the toxicity of generated outputs, such as insults and offensive language. For responsible deployment, it is crucial to rigorously safeguard these models against producing any harmful or unsafe content.
In this work, we demonstrate that existing safety mechanisms fail to prevent the generation of offensive, discriminatory, or hateful language when VLMs encounter biased or misleading contexts in text continuation tasks. In such scenarios, they often produce continuations containing harmful expressions, exposing a critical vulnerability.
@@ 1 Introduction
Figure 1: Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as Stable Diffusion.
As multimodal models, the generated content of VLMs can be significantly influenced by visual modalities, despite having the same text input. Identifying which images can effectively serve as triggers for toxic continuations remains an open challenge. Building on the substantial progress made in white-box jailbreaks Qi et al. (2024); Wang et al. (2024a), one potential approach for triggering toxic continuation is the use of adversarial images. However, optimization-based adversarial images require full access to VLM parameters, which is impractical in real-world applications, especially in VLM-as-a-Service scenarios where the VLM operates as a black-box system. Additionally, such adversarial images often exhibit noticeable distortions, making them susceptible to detection and defense mechanisms Nie et al. (2022); Zhang et al. (2023); Zhao et al. (2024). Moreover, studies have demonstrated that adversarial images typically show limited transferability between VLMs, even when crafted to target multiple VLMs simultaneously Schaeffer et al. .
These limitations motivate us to explore alternative black-box attack strategies that do not rely on full model access and generate less detectable perturbations. Current black-box methods typically rely on manually crafted pipelines to generate image-text pairs, often involving typographic attacks and the use of pre-trained diffusion models to generate query-relevant images Gong et al. (2023); Liu et al. (2023); Teng et al. (2024). However, these approaches suffer from several drawbacks. First, structure-based attacks result in a restricted attack surface, producing predictable patterns that can be easily mitigated by defense mechanisms Wang et al. (2024c). Second, such images are primarily designed to bypass sensitive word detection and are not well-suited for triggering harmful continuations in our task. Recently, automated red-teaming approaches for VLMs have gained attention. IDEATOR Wang et al. (2024b) introduced the concept of “VLMs against VLMs”, employing a red team VLM to generate jailbreak text prompts paired with images created using pre-trained diffusion models Rombach et al. (2022). While this represents significant progress, it raises critical questions about the effectiveness of pre-trained diffusion models in jailbreaking tasks. Are these pre-trained models truly optimal for inducing harmful content in VLMs? Can images generated directly by diffusers reliably trigger such behavior? To answer these questions, we explore the potential of reinforced diffusers, which leverage reinforcement fine-tuning to optimize red team image generation for more effective and targeted VLM jailbreaking.
In this paper, we propose Red Team Diffuser (RTD), a novel diffusion model framework designed to generate unrestricted images that effectively bypass safety mechanisms in VLMs. Our approach leverages a two-stage process to systematically exploit vulnerabilities in VLMs. First, we utilize an LLM to perform a greedy search for image prompts that maximize the toxicity of VLM responses, identifying highly effective image prompts for red teaming purposes. Second, we fine-tune a diffusion model through reinforcement learning, jointly optimizing a toxicity reward to amplify the harmfulness of generated outputs and an image-prompt alignment reward to ensure consistency between prompts and images as a regularization term. Figure 1 illustrates the effectiveness of our proposed method in attacking LLaVA Liu et al. (2024b), showcasing its ability to systematically induce harmful outputs. This framework reveals critical vulnerabilities in VLMs, underscoring the need for further research to address these safety challenges.
The main contributions of our work can be summarized as:
  • We introduce Red Team Diffuser (RTD), a novel framework that, to the best of our knowledge, is the first to investigate the use of unrestricted images to effectively induce harmful outputs from VLMs.
  • We propose a novel two-stage black-box attack strategy, First, an LLM performs a greedy search to identify image prompts that maximize the toxicity of VLM responses. Second, a diffusion model is fine-tuned through reinforcement learning, leveraging toxicity and alignment rewards to enhance harmful outputs.
@@
  • Experiments demonstrate that RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set. Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, as well as toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s Llama.
2 Related Work
2.1 Large Vision-Language Models
Large Vision-Language Models (VLMs) extend the functionality of Large Language Models (LLMs) by integrating visual modalities via an image encoder and modality alignment mechanisms. For example, MiniGPT-4 Zhu et al. (2023) aligned a frozen vision encoder Dosovitskiy et al. (2020) with a frozen Vicuna Chiang et al. (2023) using a single linear projection layer. InstructBLIP Dai et al. (2023), based on BLIP-2 Li et al. (2023b), introduced an instruction-tuning paradigm that leverages additional instructions to extract more useful visual features. LLaVA Liu et al. (2024b) used the text-only GPT-4 Achiam et al. (2023) to generate multimodal text-image instruction-following data and served as an end-to-end trained model that connects a vision encoder Radford et al. (2021) with LLaMA Touvron et al. (2023). While these models demonstrate significant potential in multimodal tasks, the integration of an additional visual modality introduces new avenues for attacks, distinct from those in text-only LLMs Shayegani et al. (2023b); Liu et al. (2024a); Li et al. (2023a); Ma et al. (2025).
@@ 2.1 Large Vision-Language Models
Figure 2: Overview of the proposed method. Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation. The diffusion model generates an image based on this prompt. Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
2.2 Jailbreak Attacks against VLMs
@@
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks Bagdasaryan et al. (2023); Bailey et al. (2023); Shayegani et al. (2023a); Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024); Wang et al. (2024a) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, making them less practical and more easily detectable Nie et al. (2022); Zhang et al. (2023). In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR. For example, FigStep Gong et al. (2023) converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP Liu et al. (2023); Ma et al. (2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems. To address these limitations, the concept of red team VLMs has advanced with IDEATOR Wang et al. (2024b), which autonomously generates jailbreak image-text pairs in a black-box manner. By providing a scalable and efficient solution for creating adversarial inputs, IDEATOR has demonstrated the potential of red teaming to expose vulnerabilities in VLMs. However, while IDEATOR focuses on leveraging red team VLMs, our work broadens the scope of red teaming by exploring the feasibility of red team diffusers, which are orthogonal to red team VLMs and pose greater challenges. Specifically, we investigate the capability of unrestricted images to trigger harmful content generation, highlighting the significant threats posed by such inputs to VLMs. Compared to existing approaches, our work presents a more challenging scenario by addressing the limitations of typography-based attacks and provides a rigorous framework for evaluating the robustness of VLMs against adversarial inputs.
3 Preliminary
@@
Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm Black et al. (2023) to fine-tune a pretrained diffusion model Rombach et al. (2022). Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning (RL) to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP). At each step, actions (denoising predictions) are chosen to maximize the cumulative reward. The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps. Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$. The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process. The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$. The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$. That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
@@
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator Kakade and Langford (2002):
@@
4 Methodology
4.1 Threat Model

### v2

# Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning

@@
Abstract.
The growing deployment of large Vision-Language Models (VLMs) exposes critical safety gaps in their alignment mechanisms. While existing jailbreak studies primarily focus on VLMs’ susceptibility to harmful instructions, we reveal a fundamental yet overlooked vulnerability: toxic text continuation – where VLMs produce highly toxic completions when prompted with harmful text prefixes paired with semantically adversarial images. To systematically study this threat, we propose Red Team Diffuser (RTD), the first red teaming diffusion model that coordinates adversarial image generation and toxic continuation through reinforcement learning. Our key innovations include dynamic cross-modal attack and stealth-aware optimization. For toxic text prefixes from an LLM safety benchmark, we conduct greedy search to identify optimal image prompts that maximally induce toxic completions. The discovered image prompts then drive RL-based diffusion model fine-tuning, producing semantically-aligned adversarial images that boost toxicity rates. Stealth-aware optimization introduces joint adversarial rewards that balance toxicity maximization (via Detoxify classifier) and stealthiness (via BERTScore), circumventing traditional noise-based adversarial patterns. Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% over text-only baselines on the original attack set and 8.91% on unseen set, proving generalization capability. Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA. Our findings expose two critical flaws in current VLM alignment: (1) failure to prevent toxic continuation from harmful prefixes, and (2) overlooking cross-modal attack vectors. These results necessitate a paradigm shift toward multimodal red teaming in safety evaluations. We will publicly release the RTD framework to facilitate future research on robust VLM alignment.
@@ Abstract.
Red Teaming, Vision-Language Models, Diffusion Models, Reinforcement Learning
ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Figure 1. Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as general-purpose diffusion model (Rombach et al., 2022).
The rapid deployment of large Vision-Language Models (VLMs), such as GPT-4o (Achiam et al., 2023) and Gemini (Team et al., 2023), has exposed critical blind spots in their multimodal alignment mechanisms. While existing safety strategies predominantly focus on preventing responses to explicit harmful instructions (e.g., ”How to build a bomb?”) through techniques like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), they often fail to address the risks posed by toxic text continuation tasks - where VLMs amplify harmful content when completing partially toxic text prompts paired with semantically adversarial images (Figure 1). This vulnerability originates from two research gaps: (a) Current red teaming methods focus on explicit instruction attacks rather than contextual toxicity induction. (b) Existing safety protocols neglect cross-modal attack vectors where images amplify output toxicity.
The multimodal nature of VLMs introduces unique safety challenges: identical text inputs paired with different images can yield drastically different toxicity levels (Liu et al., 2023), sometimes in unintuitive ways. For instance, Li et al. (Li et al., 2024) demonstrated that pairing jailbreak prompts with even a blank image can significantly increase attack success rates. However, systematically identifying images that effectively trigger toxic continuations remains an open and challenging problem. Recent progress in white-box jailbreak techniques (Qi et al., 2024; Wang et al., 2024b) has explored the use of adversarial images optimized directly against model parameters. While powerful, such gradient-based approaches are impractical in real-world scenarios, particularly within black-box VLM-as-a-Service contexts where full model access is unavailable. Additionally, adversarial images generated through gradient-based methods typically exhibit noticeable distortions, making them susceptible to detection and defense mechanisms (Nie et al., 2022; Zhang et al., 2023; Zhao et al., 2024). Moreover, studies have demonstrated that these adversarial images typically show limited transferability between VLMs, even after successfully jailbreaking multiple VLMs simultaneously (Schaeffer et al., [n. d.]).
These limitations motivate us to develop stealthy black-box attacks that produce natural adversarial images without full model access. Existing black-box jailbreak methods typically adopt suboptimal paradigms, rely on manually engineered pipelines involving typographic manipulations or pre-trained diffusion models to generate query-relevant images (Gong et al., 2023; Liu et al., 2023; Teng et al., 2024). However, such methods have critical drawbacks. First, structure-based attacks inherently constrain the attack surface, resulting in predictable patterns that are easily mitigated by modern defense mechanisms (Wang et al., 2024a) and detectable by OCR systems. Second, such images are primarily designed to bypass sensitive text detection, rather than actively inducing harmful continuations, limiting their applicability to implicit toxicity tasks.
More recently, automated red-teaming frameworks have emerged to systematically uncover vulnerabilities in VLMs. For instance, IDEATOR (Wang et al., 2024c) introduced the ”VLMs against VLMs” paradigm, employing a red-team VLM to generate jailbreak prompts paired with images from general-purpose diffusion models (Rombach et al., 2022). While innovative, this approach inherits a fundamental limitation: the inherent conflict between diffusion models’ original training objective (faithful visual-textual alignment) and red teaming goals (toxicity maximization). Off-the-shelf diffusers like Stable Diffusion (Rombach et al., 2022) are optimized to avoid generating harmful content during pretraining, creating an intrinsic bias against jailbreaking effectiveness.
This highlights a critical yet unexplored research opportunity: Can we reconfigure diffusion models to actively induce toxicity while maintaining semantic plausibility? Current approaches treat image generation and toxicity induction as decoupled processes - first generating ”safe” images with pretrained diffusers, then hoping they coincidentally trigger harmful outputs. Our work overturns this passive paradigm through reinforced diffusion alignment, where the diffusion model is actively tuned to: (1) maximize toxicity through RL-driven reward signals; (2) preserve semantic coherence with image prompts; (3) dynamically adapt to target VLM’s defense patterns. By transforming diffusion models from general-purpose generators into active toxicity amplifiers, we establish a new paradigm for multimodal red teaming that fundamentally rethinks how visual modalities can be weaponized against VLMs.
In this work, we propose Red Team Diffuser (RTD), a novel framework that redefines multimodal jailbreak through semantically adversarial image synthesis. Departing from conventional static attack patterns, RTD introduces a dynamic two-phase optimization paradigm. First, we leverage the contextual reasoning capabilities of an LLM to perform a greedy search over image prompts, identifying those that maximize the toxicity of VLM responses and are thus highly effective for red teaming. Second, the discovered prompts drive RL-based fine-tuning of a diffusion model, using a dual-objective reward: a toxicity reward to amplify the harmfulness of the induced outputs, and an image-prompt alignment reward to ensure semantic consistency as a regularization term. This transforms standard diffusion models into toxicity-aware generators, fundamentally altering their output distribution toward vulnerable regions of target VLMs’ multimodal alignment space.
The main contributions of our work are threefold:
  • New Attack Paradigm: We propose Red Team Diffuser (RTD), the first framework to integrate reinforcement learning with diffusion models for semantically adversarial image generation, establishing a new paradigm for multimodal red teaming. Unlike prior approaches that rely on static patterns or unmodified diffusers, RTD actively co-optimizes both image prompts and diffusers for toxicity induction.
  • Technical Innovations: We employ an LLM to automatically generate diverse image prompts, followed by iterative toxicity evaluation to identify the most effective triggers. Furthermore, we introduce an RL-based fine-tuning strategy with dual rewards, balancing toxicity maximization and semantic preservation, thereby ensuring both high attack potency and natural image quality.
@@
  • Empirical Contributions: We demonstrate a 10.69% increase in toxicity rate over text-only attacks on LLaVA. Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, along with toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s LLaMA. Our findings reveal a critical vulnerability in cross-modal toxicity amplification, where semantically adversarial images exacerbate text-based harm.
2. Related Work
2.1. Large Vision-Language Models
Large Vision-Language Models (VLMs) extend Large Language Models (LLMs) by integrating visual modalities through an image encoder and a cross-modal alignment module. This design enables VLMs to perform instruction-following and open-ended reasoning over both text and images. Recent open-source VLMs have explored various architecture designs. LLaVA (Liu et al., 2024a) connects a CLIP encoder (Radford et al., 2021) with LLaMA (Touvron et al., 2023), using GPT-4 (Achiam et al., 2023) to synthesize multimodal instruction-following data for end-to-end fine-tuning. MiniGPT-4 (Zhu et al., 2023) employs a linear projection to align a frozen ViT (Dosovitskiy et al., 2020) with Vicuna (Chiang et al., 2023), using pretraining on large-scale image-text pairs followed by high-quality fine-tuning. InstructBLIP (Dai et al., 2023), based on BLIP-2 (Li et al., 2023b), introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding. While these models achieve impressive performance, the incorporation of visual inputs significantly broadens the attack surface compared to text-only LLMs. Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior (Shayegani et al., 2023b; Liu et al., 2024b; Li et al., 2023a; Ma et al., 2025). These findings highlight the urgent need to evaluate VLM robustness under multimodal threats—an area this work directly addresses.
@@ 2.1. Large Vision-Language Models
Figure 2. Overview of the proposed method. Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation. The diffusion model generates an image based on this prompt. Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
2.2. Jailbreak Attacks against VLMs
@@
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks (Bagdasaryan et al., 2023; Bailey et al., 2023; Shayegani et al., 2023a; Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024; Wang et al., 2024b) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, making them less practical and more easily detectable (Nie et al., 2022; Zhang et al., 2023). In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR. For example, FigStep (Gong et al., 2023) converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP (Liu et al., 2023; Ma et al., 2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems. To address these limitations, the concept of red team VLMs has progressed with IDEATOR (Wang et al., 2024c), which autonomously generates jailbreak image-text pairs in a black-box setting. By offering a scalable and automated pipeline for discovering jailbreak prompts, IDEATOR demonstrates the promise of using VLMs themselves to expose alignment weaknesses. While IDEATOR focuses on prompt-level red teaming, our work introduces a fundamentally different and complementary attack paradigm: reinforced diffusion-based image optimization. Rather than relying on VLM-generated inputs, we investigate how semantically adversarial images—crafted via reinforcement learning—can effectively induce toxic text continuations when paired with fixed text prefixes. This setup presents a significantly more challenging threat model: it removes reliance on typographic or prompt engineering tricks and explores the vulnerability of VLMs to unrestricted, natural-looking visual stimuli. By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks, highlighting a critical blind spot in existing alignment strategies.
3. Preliminary
@@
Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm (Black et al., 2023) to fine-tune a pretrained diffusion model (Rombach et al., 2022). Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP). At each step, actions (denoising predictions) are chosen to maximize the cumulative reward. The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps. Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$. The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process. The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$. The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$. That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
@@
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator (Kakade and Langford, 2002):
@@
4. Methodology
4.1. Threat Model